OpenStudy (anonymous):

Suppose that the number of cars manufactured at an automobile plant varies jointly as the number of workers and the time they work. If 220 workers can produce 198 cars in 5 hours, find the number of cars that 175 workers can produce in 6 hours.

OpenStudy (anonymous):

In one hour, one worker can produce: \[\frac{198}{220\times 5} = \frac{9}{50} \text{ cars} \] \[\implies \frac{9}{50} \times 175 \times 6\] Can be produced in 6 hours by 175.